OpenStudy (anonymous):

help please? In solving the equation (x + 4)(x – 7) = -18, Eric stated that the solution would be x + 4 = -18 => x = -22 or (x – 7) = -18 => x = -11 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

the reason eric's solution doesn't work is cebause he is not solving at y= 0, he is solving at y=-18. in order to find the roots of an equation, you must solve it at y=0 so in this case we must first multiply out our equation: (x+4)(x-7)=x^2 -3x-28 = -18 to get y= 0 our equation changes x^2-3x-28 = -18 x^2-3x -10=0 (by adding 18 to both sides) when we factor this equation we get: x^2-3x -10= (x-5)(x-2)=0 so the solutions to the equation are actually: x-5=0 x=5 x+2=0 x=-2 when we check these in our original equation they should work

OpenStudy (anonymous):

thank you thank you thank you! that one stumped me so much!

OpenStudy (anonymous):

np